### 05 Mar 2026 v3

@@ VI.2 Junction conditions:
    Testing the denominator of the last equation for zero leads us to:
@@
In Section II, a new string fluid solution was introduced. There, it was presented under the assumption that it extends over the entire spacetime, $r\in[0,\infty)$. In the present section, we analyze instead the case in which the solution of Section II, in the presence of a negative cosmological constant, is defined only in the interval $r\in[0,h]$, i.e., it can represent the interior region of the Reissner–Nordström solution, as depicted in Scheme 2. It is worth mentioning that, in Section VI, we have already highlighted the main features of the Reissner–Nordström solution that are relevant to this work.
VII.1 Representation of the New Fluid of Strings solution as an Interior Region with an Integrable Singularity
@@
Thus, the trace of the effective interior energy–momentum tensor with a cosmological constant behaves as $R\sim r^{-2}$ at short length scales. This behavior implies that the equations of motion (28) are integrable in the vicinity of the origin and remain free of pathological singularities. Consequently, condition IV 1 is satisfied, allowing this new string fluid model to consistently represent the interior region as an integrable singularity.
The solution is given by:
@@
In order to test condition IV 2 for the interior region to be represented as an integrable singularity, we note that $f(r)$ is an increasing function, starting from its minimum value $f(r=0)=1-\frac{2M}{b}<0$, with $\frac{2M}{b}>1$, and, as we shall see below, reaching $f(h)=0$. Indeed, $df/dr=\frac{2M}{r^{2}}\left(1-\exp(-r/b)\left(1+\frac{r}{b}\right)\right)+\frac{2r}{l^{2}}>0$ for $r\in[0,h]$.
@@ VIII Discussion and Summary
In Section II, we present a new fluid-of-strings (fs) black hole solution defined over the entire domain $r\in[0,\infty)$ and characterized by the equation of state $\rho^{(fs)}(r)$ and $p^{(fs)}(r)$. Following Ref. Aoki et al. (2021), where the integral $\int_{0}^{\infty}4\pi r^{2}T^{0}{}_{0},dr$ is interpreted as the conserved energy associated with a timelike Killing vector, we note that this integral diverges for the standard string-cloud energy density. Motivated by this behavior, we introduce a modified energy density for the fluid of strings, which can be interpreted as a geometrical screening of the original string-cloud distribution. As a result, the aforementioned integral converges to a finite value equal to $M$, which can be interpreted as the conserved energy charge. Although the Ricci scalar diverges near the origin as $R\sim r^{-2}$, these divergences are integrable, and the spacetime remains free of curvature singularities, allowing the solution to be classified as an integrable singularity. Finally, we show that the solution is free of a potentially unstable de Sitter core and of an inner horizon, whose presence is associated with a breakdown of causality.
Motivated by the fact that several black hole solutions sourced by matter fields leading to a destructive central singularity have attracted considerable interest in the literature, and in analogy with Ref. Ovalle (2024), we have addressed, as a second problem, the existence of an internal geometry inside the event horizon of a generic black hole spacetime. Within this framework, the external geometry is kept completely general, with the Schwarzschild solution appearing only as a particular case. We have focused on scenarios in which the internal structure is endowed with an integrable singularity, in order to avoid the emergence of a potentially unstable de Sitter core and the presence of an inner horizon, since predictability breaks down after crossing the inner horizon. We have investigated the nature of the matter sources capable of supporting such an internal geometry with an integrable singularity, specifically those associated with clouds of strings and with the new string fluid introduced in this work. To this end, in Section IV we have defined the conditions that the internal geometry must satisfy in order to represent an integrable singularity. In Section V we have established the junction conditions between the internal and generic external black hole geometries. We have shown that the second fundamental form of the Israel–Darmois junction conditions Israel (1966) leads to the continuity of the temperature across the interface between the internal and external regions. Furthermore, from the analysis of the junction conditions, we have determined that a discontinuity in the tangential pressure implies the existence of a phase transition at the location of the event horizon, and vice versa.
@@
In Sections VI and VI, we studied the cases in which the internal geometry of the Reissner–Nordström spacetime is described, respectively, by a string cloud and by our newly proposed string fluid geometry. In both scenarios, the parameters of the corresponding internal geometries satisfy the conditions listed in Section IV, ensuring the presence of an integrable singularity near the origin, the absence of a potentially unstable de Sitter core, and the absence of an internal horizon that could lead to a breakdown of causality. In Section VI, we determined the allowed values of the energy density parameter $a$ and of the internal cosmological constant required to satisfy both Israel junction conditions at the interface between the external RN region and the internal string cloud region. In addition, we identified a critical value of the electric charge for which no phase transition occurs in the location of the interface between the two regions. In Section VI, we derived the conditions that the AdS radius, the electric charge, the screening parameter $b$, and the event horizon must satisfy in order to fulfill both Israel junction conditions at the interface between the external RN region and the internal region described by the new string fluid. Moreover, we numerically determined the critical value of the screening parameter for which no phase transition occurs in the location of the interface between the two regions.
Appendix A A brief revision of the cloud of string solution
@@
Appendix B Some general aspects of the fluid of strings solution

### 09 Mar 2026 v4

@@ VI.2 Junction conditions:
    Testing the numerator of the last equation for zero leads us to:
@@
In Section II, a new string fluid model was introduced. There, it was presented under the assumption that it extends over the entire spacetime, $r\in[0,\infty)$. In the present section, we analyze instead the case in which the model of Section II, in the presence of a negative cosmological constant, is defined only in the interval $r\in[0,h]$, i.e., it can represent the interior region of the Reissner–Nordström solution, as depicted in Scheme 2. It is worth mentioning that, in Section VI, we have already highlighted the main features of the Reissner–Nordström solution that are relevant to this work.
VII.1 Representation of the New Fluid of Strings model as an Interior Region with an Integrable Singularity
@@
Thus, the trace of the effective interior energy–momentum tensor with a cosmological constant behaves as $R\sim r^{-2}$ at short length scales. This behavior implies that the equations of motion (14) are integrable in the vicinity of the origin and remain free of pathological singularities. Consequently, condition IV 1 is satisfied, allowing this new string fluid model to consistently represent the interior region as an integrable singularity. The geometry is given by:
@@
In order to test condition IV 2 for the interior region to be represented as an integrable singularity without an inner horizon, we note that $f(r)$ is an increasing function, starting from its minimum value $f(r=0)=1-\frac{2M}{b}<0$, with $\frac{2M}{b}>1$, and, as we shall see below, reaching $f(h)=0$. Indeed, $df/dr=\frac{2M}{r^{2}}\left(1-\exp(-r/b)\left(1+\frac{r}{b}\right)\right)+\frac{2r}{l^{2}}>0$ for $r\in[0,h]$.
@@ VIII Discussion and Summary
In this work, motivated by the existence of an inner horizon and a non-integrable singularity in the Reissner–Nordström solution (RN), we study the physical and thermodynamic properties of a scenario in which the RN black hole possesses an interior region with an integrable singularity and no inner horizon, instead of a pointlike mass. We have investigated the possibility that the nature of the matter sources capable of supporting such an internal geometry with an integrable singularity (IS) corresponds to clouds of strings (CS) and fluids of strings (FS). In this connection, motivated by the fact that the conserved charge energy in the CS case diverges, we have proposed a model for the internal region based on an FS in which the energy density, which can be interpreted as a geometrical screening of the original CS distribution, leads to a finite conserved charge energy. We have shown that both the CS and FS models used in this work can represent an interior region with an IS and that they lack an inner horizon where predictability would be lost.
To carry out this work, in Section IV we have determined the conditions that a general internal geometry must satisfy in order to represent an integrable singularity. In Section V we have defined the Israel–Darmois Israel (1966) junction conditions between a general internal region and a general external BH geometry. In this context, we have shown that satisfying the junction conditions requires the continuity of temperature at the interface, while discontinuities in the tangential pressure signal the presence of phase transitions. It is worth mentioning that these general conditions are also of interest for future studies of other BH exterior regions that possess a central singularity instead of the RN one, for example BHs with cosmic void density profiles Lustosa et al. (2025), BHs surrounded by dark matter Xu et al. (2018), hairy BHs Ovalle et al. (2021), and quintessential BHs Kiselev (2003); Estrada et al. (2025).
@@
Using the framework developed in the previous paragraph, in Sections VI and VII we studied the cases in which the internal geometry of the Reissner–Nordström spacetime is described, respectively, by a string cloud and by our newly proposed string fluid model. In both scenarios, the parameters of the corresponding internal geometries satisfy the conditions listed in Section IV, ensuring the presence of an integrable singularity near the origin, the absence of a potentially unstable de Sitter core, and the absence of an internal horizon that could lead to a breakdown of causality.
In Section VI, we determined the allowed values of the energy density parameter $a$ and of the internal cosmological constant required to satisfy both Israel junction conditions at the interface between the external RN region and the internal string cloud region. In addition, we identified a critical value of the electric charge for which no phase transition occurs in the location of the interface between the two regions.
In Section VII, we derived the conditions that the AdS radius, the electric charge, the screening parameter $b$, and the event horizon must satisfy in order to fulfill both Israel junction conditions at the interface between the external RN region and the internal region described by the new string fluid. Moreover, we numerically determined the critical value of the screening parameter for which no phase transition occurs in the location of the interface between the two regions.
Appendix A A brief revision of the cloud of string model
@@
Appendix B Some general aspects of the fluid of strings model